Solve the system of equations. Express your answer in fractions. 5x – 7y = 31 4x + 3y = 19
Answer #1

The given system of eanations are 56-79=31 - 44 of 3y = 19 6 6. from 5x = 31t ty P, xo 3147g. 5 put x = 3479 in @ we have 4.1